OpenStudy (anonymous):

what equation represents the line that runs throught the points (1,5) and has a slope of -2?

OpenStudy (anonymous):

try to put it in point slope form: y – y1 = m(x – x1)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &({\color{red}{ 1}}\quad ,&{\color{blue}{ 5}})\quad \end{array} \\\quad \\ slope = {\color{green}{ m}}= -2 \\ \quad \\ y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and solve for "y"}\\ \qquad \uparrow\\ \textit{point-slope form} \)